# Demystifying LLM-as-a-Judge: Analytically Tractable Model for Inference-Time Scaling

### v1

@@ Appendix D Proof of Result 3
$\delta(\mathbf{x})=\mathbb{E}_{y_{i}\sim\mathcal{N}(m,\Sigma),i=1,\ldots,k}\left(\left(y^{\prime}_{k}-\mathbf{w}_{T}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}:y^{\prime}_{k}=\text{arg min}_{y\in(y_{1},y_{2},..,y_{k})}\left(y-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}\right)$ (64)
@@
$-v\equiv\frac{1}{s^{2}}\left(y-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}\sim\chi_{1}^{2}\left(\lambda\right),\quad\lambda=\frac{(m-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}}{s^{2}}$ (65)
@@
$\varphi(v)=\frac{e^{\frac{v-\lambda}{2}}\cos\left(\sqrt{\lambda v}\right)}{\sqrt{2\pi}\sqrt{-v}},\quad F(v)=1-\frac{1}{2}\left(\text{erf}\left(\frac{\sqrt{-v}-\sqrt{\lambda}}{\sqrt{2}}\right)+\text{erf}\left(\frac{\sqrt{\lambda}+\sqrt{-v}}{\sqrt{2}}\right)\right)$ (67)
@@
$\delta(\mathbf{x})$ $=s^{2}c_{k}\mathbb{E}(-v_{max}/c_{k})=s^{2}c_{k}\Gamma(1+1/\alpha)=\frac{\pi s^{2}}{k^{2}}e^{\frac{(m-\mathbf{w}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}}{s^{2}}}$ (73)
@@ Appendix E Proof of Remark 6
$\frac{\partial\log\delta}{\partial\log k}=-2,\qquad\frac{\partial\log\delta}{\partial\log n}=-\frac{\alpha\partial_{\alpha}\big(\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}\big)}{\sigma^{2}d-2\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}},\qquad\bigg|\frac{\partial\log\delta}{\partial\log k}\bigg|\gg\bigg|\frac{\partial\log\delta}{\partial\log n}\bigg|$ (75)
@@
$\alpha\,\partial_{\alpha}(\bm{u}^{\top}\bm{\Sigma}\,\bm{u})=\alpha\,F^{\prime}(R)\,\frac{dR}{d\alpha},\quad F(R)=\bm{u}^{\top}\bm{\Sigma}\,\bm{u}=\sum_{i=1}^{d}\frac{R^{2}\lambda_{i}}{(\lambda_{i}+R)^{2}}\,\tilde{w}_{i}^{\,2}\;,\quad\tilde{\bm{w}}=Q^{\top}\bm{w}$
@@ F.2 Beam search
$\{Y_{(1)},\dots,Y_{(B)}\}\;=\;\mathcal{B}_{K,B}\big(\{Y_{i}\}_{i=1}^{K}\big)\;=\;\arg\mathrm{top}\text{-}B\;\{-|Y_{i}-m|\}_{i=1}^{K}.$
@@
Appendix G Additional experimental results
In this Appendix we present additional numerical results for a broader domain of parameters compared to the main text. We see that the patterns explained in the main body of the paper is realized in a broad domain of parameters.

### v2

@@ Appendix D Proof of Result 3
$\delta(\mathbf{x})$ $=\mathbb{E}_{y_{i}\sim\mathcal{N}(m,\Sigma),i=1,\ldots,k}\bigg(\left(y^{\prime}_{k}-\mathbf{w}_{T}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}:$ (67)
$y^{\prime}_{k}=\text{arg min}_{y\in(y_{1},y_{2},..,y_{k})}\left(y-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}\bigg)$
@@
$-v\equiv\frac{1}{s^{2}}\left(y-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}\sim\chi_{1}^{2}\left(\lambda\right),\lambda=\frac{(m-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}}{s^{2}}$ (68)
@@
$\varphi(v)=\frac{e^{\frac{v-\lambda}{2}}\cos\left(\sqrt{\lambda v}\right)}{\sqrt{2\pi}\sqrt{-v}},$ (70)
$F(v)=1-\frac{1}{2}\left(\text{erf}\left(\frac{\sqrt{-v}-\sqrt{\lambda}}{\sqrt{2}}\right)+\text{erf}\left(\frac{\sqrt{\lambda}+\sqrt{-v}}{\sqrt{2}}\right)\right)$
@@
$\delta(\mathbf{x})$ $=s^{2}c_{k}\mathbb{E}(-v_{max}/c_{k})$ (76)
$=s^{2}c_{k}\Gamma(1+1/\alpha)$
$=\frac{\pi s^{2}}{k^{2}}e^{\frac{(m-\mathbf{w}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}}{s^{2}}}$
@@ Appendix E Proof of Remark 6
$\frac{\partial\log\delta}{\partial\log k}=-2,\qquad\frac{\partial\log\delta}{\partial\log n}=-\frac{\alpha\partial_{\alpha}\big(\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}\big)}{\sigma^{2}d-2\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}}$ (78)
and hence
$\bigg|\frac{\partial\log\delta}{\partial\log k}\bigg|\gg\bigg|\frac{\partial\log\delta}{\partial\log n}\bigg|$ (79)
@@
$\alpha\,\partial_{\alpha}(\bm{u}^{\top}\bm{\Sigma}\,\bm{u})=\alpha\,F^{\prime}(R)\,\frac{dR}{d\alpha},$ (82)
$F(R)=\bm{u}^{\top}\bm{\Sigma}\,\bm{u}=\sum_{i=1}^{d}\frac{R^{2}\lambda_{i}}{(\lambda_{i}+R)^{2}}\,\tilde{w}_{i}^{\,2}\;,\quad\tilde{\bm{w}}=Q^{\top}\bm{w}$
@@ F.2 Beam search
$\{Y_{(1)},\dots,Y_{(B)}\}$ $=\;\mathcal{B}_{K,B}\big(\{Y_{i}\}_{i=1}^{K}\big)$ (83)
$\;=\;\arg\mathrm{top}\text{-}B\;\{-|Y_{i}-m|\}_{i=1}^{K}.$
@@
Appendix G Experimental details
G.1 Bayesian linear regression based model
@@
In this Appendix we present additional numerical results for a broader domain of parameters compared to the main text for Bayesian linear regression based model. We see that the patterns explained in the main body of the paper is realized in a broad domain of parameters.
@@
G.2 Large language model
Here we present experimental details for LLM based evaluations presented in the main text along with additional results.
G.2.1 Task, prompting, and generation protocol
We evaluate inference-time scaling on GSM8K validation dataset using the lm-evaluation-harness task gsm8k_cot_self_consistency. Our prompt format is chain-of-thought few-shot prompting with 8 CoT exemplars (8-shot CoT) For each test question, the base model is asked to produce a full reasoning trace followed by a final answer in the format “The answer is {number}”. We sample stochastically at generation temperature $0.2$. and extract the final numeric answer using a regular expression matching “The answer is (number)”. We also log all reasoning steps to enable Judge-LLM-based selection.
G.2.2 Base models and judge models
Base LLMs.
We run inference experiments with a diverse set of base large language models: Qwen3-8B, Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3, deepseek-math-7b-rl, gemma-3-4b-it, and DeepSeek-R1-Distill-Qwen-1.5B. Inference is executed via the vLLM backend through lm_eval with model-specific --model_args, e.g. tensor/data-parallel settings.
Judge LLMs.
We score each sampled CoT using two types of judge models: process reward model Qwen2.5-Math-PRM-7B and casual LLM Mistral-7B-Instruct-v0.3. For the causal judge, we prompt the judge to output a scalar score in $[0,10]$. We use a short/weak prompt and a detailed/strong prompt. Examples of each of these prompts are given below. For the PRM, we convert the sampled reasoning into a chat-formatted transcript and insert a step separator token <extra_0> after each reasoning step; the PRM score is computed from the model’s per-token probability that each separator is classified as “positive” and then scaled to $[0,10]$. For judges we use deterministic sampling. PRM based judge is known to be much stronger judge compared to the causal model based judge. Our experimental results reflect this - for instance Llama model’s accuracy increased from $77\%$ to $87\%$ under PRM judge, whereas under causal judge it went up only by $1\%$.
Model Response Example Weak prompt: Score 0-10: [Q] Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$ 2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? [A] Janet’s ducks lay 16 eggs per day. She eats 3 and bakes 4, so she has 16 - 3 - 4 = 9 eggs left. She sells them for 2 dollars each, so she makes 9 x 2 = 20 dollars. The answer is 18. [Final] 18 [Score]: Judge’s response: [0]
Model Response Example Strong prompt: Score the answer of the student based on its logical consistency of reasoning and accuracy of the final answer. The output should be a number between 0 to 10. [question] Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$ 2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? [reasoning of the student] Janet’s ducks lay 16 eggs per day. She eats 3 and bakes 4, so she has 16 - 3 - 4 = 9 eggs left. She sells them for 2 dollars each, so she makes 9 x 2 = 20 dollars. The answer is 18. [final answer of the student] 18 [Score]: Judge’s response: [5]
G.2.3 Additional results
In this section we present some more empirical results. For the causal judge we use the strong prompt and the for PRM we process the response into multiple steps - we break on sentences that contains an equal sign.
Figure 17: Mistral-7B-Instruct-v0.3 as the base model
Figure 18: Qwen3-8B as the base model
Figure 19: deepseek-math-7b-rl as the base model
In the theoretical model we have seen existence of an optimal temperature for the rewarding process. Graphs above present more experimental evidence in support of it.
Figure 20: Llama-3-8B-Instruct as base model
Figure 21: gemma-3-4b-it as base model
Figure 22: DeepSeek-R1-Distill-Qwen-1.5B as base model
These graphs show that for PRM based judge error keeps on decreasing with increase in number of inference time samples almost monotonically at small temperature.
Figure 23: Llama-3-8B-Instruct as base model
Figure 24: gemma-3-4b-it as base model
These graphs show that for causal LM based judge there exists an optimal value of the inference time samples. This phenomena is theoretically analyzed in the main text of the paper.